# Graph-based Retrieval Augmented Generation for Dynamic Few-shot Text Classification

### v1

@@ 4.4. Case Study
4.5. Discussion
According to the aforementioned experimental results, we have identified several key characteristics of NN-based models and RAG-based models.
On the one hand, despite the rapid development of LLMs in recent years, they can still yield sub-optimal classification accuracy in DFSTC tasks due to the lengthy LLM inputs. When there is a sufficient amount of labeled data available in each dynamic updation round, the NN-based model Entailment can also achieve comparable or even better performance than traditional RAG based models other than GORAG.
On the other hand, compared with NN-based models, benefited from the time efficient indexing of the side information and the rich pre-trained knowledge within the LLM, by filter out some unrelated labels and shorten the LLM inputs, RAG-based models are better suited for DFSTC task when the target labels are updated frequently and there only existes a limit number or even none of the labeled data is avaiable for each dynamic updation round.
5. Conclusion
In this paper, we propose GORAG, a Graph-based Online Retrieval Augmented Generation framework for the Dynamic Few-shot Text Classification (DFSTC) task. Extensive experiments on two text datasets with different characteristics demonstrate the effectiveness of GORAG in classifying texts with only a limited number or even no labeled data. Additionally, GORAG shows its effectiveness in adapting to the dynamic updates of target labels by retrieving candidate labels to filter the large target label set in each update round. With extensive of ablation studies, we confirm that GORAG’s adaptive retrieval, edge weighting, and online indexing mechanisms contribute to its effectiveness. For future work, we aim to further enhance GORAG’s performance and explore its application in more complex and diverse scenarios, as well as further improve the efficiency of GORAG’s online indexing mechanism.

### v2

@@ E.5. Case Study
Table 8. The macro recall score and hallucination rate (Hall. Rate) on WOS dataset.
Metrics Category Model Round
R1 R2 R3 R4
Recall NN-based Entailment 0.3502 0.3475 0.3908 0.3021 0.3415 0.3442 0.2327 0.2406 0.2432 0.1789 0.1792 0.1837
Long Context RAG NaiveRAG 0.3324 0.3417 0.3398 0.1615 0.1652 0.1650 0.0943 0.1013 0.0872 0.0568 0.0805 0.0628
Propositionizer 0.1032 0.1131 0.1114 0.0947 0.1210 0.1187 0.1352 0.1298 0.1254 0.1373 0.1345 0.1402
Compression RAG LongLLMLingua 0.3139 0.3105 0.3258 0.1349 0.1388 0.1350 0.0862 0.0837 0.0872 0.0616 0.0610 0.0628
GraphRAG 0.3791 0.3646 0.3732 0.1535 0.1519 0.1557 0.1009 0.1014 0.1015 0.0902 0.0940 0.0912
LightRAG 0.3462 0.3441 0.3519 0.1544 0.1507 0.1466 0.1241 0.1257 0.1226 0.1129 0.1180 0.1101
GORAG 0.4387 0.4659 0.4874 0.3074 0.3472 0.3483 0.2371 0.2513 0.2445 0.1960 0.1970 0.2013
Hal. Rate NN-based Entailment 54.53% 58.96% 51.63% 44.76% 36.62% 37.76% 32.57% 28.14% 27.35% 29.68% 26.86% 24.13%
Long Context RAG NaiveRAG 3.00% 2.32% 2.63% 6.45% 6.16% 6.28% 6.93% 7.46% 7.52% 9.96% 9.50% 9.83%
Propositionizer 6.85% 5.88% 6.97% 7.40% 6.52% 6.09% 5.97% 6.44% 6.30% 5.78% 6.11% 5.52%
Compression-RAG LongLLMLingua 6.50% 6.29% 5.63% 16.45% 16.05% 16.28% 19.63% 20.47% 20.52% 21.96% 21.35% 21.83%
GraphRAG 2.40% 2.23% 2.07% 6.03% 5.97% 6.18% 7.20% 7.58% 7.03% 6.90% 6.39% 6.49%
LightRAG 1.70% 1.24% 1.37% 2.59% 2.78% 3.21% 4.66% 4.55% 4.58% 5.89% 5.80% 5.73%
GORAG 0.66% 1.16% 0.74% 0.30% 0.68% 0.65% 0.41% 0.57% 0.68% 0.40% 0.43% 0.64%
Table 9. The macro recall score and hallucination rate (Hall. Rate) on IFS-Rel dataset.
Metrics Category Model Round
R1 R2 R3 R4
Recall NN-based Entailment 0.3191 0.6015 0.5515 0.2172 0.3043 0.3330 0.1225 0.1802 0.2081 0.1093 0.1267 0.1490
Long Context RAG NaiveRAG 0.4039 0.4194 0.4292 0.2320 0.2009 0.2115 0.1147 0.1163 0.1097 0.0871 0.0895 0.0912
Propositionizer 0.1053 0.1146 0.1104 0.1124 0.1201 0.1197 0.0824 0.1029 0.0902 0.0744 0.0717 0.0799
Compression RAG LongLLMLingua 0.4426 0.4625 0.4044 0.1879 0.1559 0.2158 0.0947 0.0906 0.1113 0.0698 0.0822 0.0696
GraphRAG 0.4264 0.4139 0.4208 0.2116 0.2008 0.2047 0.1213 0.1179 0.1185 0.0906 0.0857 0.0883
LightRAG 0.4221 0.3985 0.4382 0.2344 0.2150 0.2429 0.1270 0.1313 0.1276 0.0902 0.0926 0.0918
GORAG 0.4588 0.4681 0.5000 0.2348 0.2175 0.2405 0.1279 0.1295 0.1290 0.0950 0.0886 0.0919
Hal. Rate NN-based Entailment 8.10% 15.63% 4.69% 26.41% 19.14% 6.25% 26.56% 17.86% 6.67% 6.48% 5.90% 3.28%
Long Context RAG NaiveRAG 1.56% 4.25% 1.78% 3.05% 2.97% 3.13% 3.02% 3.70% 3.54% 4.02% 4.18% 3.63%
Propositionizer 7.82% 7.34% 11.25% 5.78% 6.40% 6.72% 4.84% 6.09% 7.03% 8.13% 8.59% 6.72%
Compression-RAG LongLLMLingua 3.59% 2.97% 6.56% 2.12% 6.23% 5.47% 6.36% 6.00% 3.50% 3.29% 5.65% 8.88%
GraphRAG 4.25% 1.56% 1.25% 2.66% 3.98% 2.89% 3.39% 3.75% 3.80% 3.91% 3.24% 3.09%
LightRAG 3.28% 2.66% 2.03% 2.73% 2.97% 2.27% 1.98% 1.56% 1.67% 2.46% 2.23% 2.30%
GORAG 0.31% 0.63% 0.16% 1.78% 2.81% 2.58% 1.30% 0.78% 1.09% 1.48% 1.02% 1.25%